# Complexity Analysis of Normalizing Constant Estimation: from Jarzynski Equality to Annealed Importance Sampling and beyond

### v1

@@ 5 Analysis of the Annealed Importance Sampling
    Show that it suffices to obtain an accurate estimate $\widehat{Z}_{0}$ and initialization distribution $\widehat{\pi}_{0}$, together with sufficiently small KL divergences $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$ and $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$, which quantify the closeness between the continuous dynamics and the discretization error in implementation, respectively.
  3. 3.
@@
    Using the chain rule, decompose $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$ into the sum of KL divergences between each pair of transition kernels $F_{\ell}$ and $F^{*}_{\ell}$ (i.e., the sum of green “distances”). As in the proof of the convergence of JE (Thm. 6), $F^{*}_{\ell}$, a transition kernel from $\pi_{\theta_{\ell-1}}$ to $\pi_{\theta_{\ell}}$, is realized by ALD with a compensatory vector field, ensuring the SDE exactly follows the trajectory $(\pi_{\theta})_{\theta\in[\theta_{\ell-1},\theta_{\ell}]}$. Similarly, by applying the chain rule and Girsanov theorem, we can express $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$ as the sum of the blue “distances”, allowing for a similar analysis.
  4. 4.
@@
    Finally, derive three necessary conditions on the time steps $\theta_{\ell}$ to control both $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$ and $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$. Choosing a proper schedule yields the desired complexity bound.
@@
Our proposed algorithm consists of two phases: first, estimating $Z_{0}$ by TI, which is provably efficient for well-conditioned distributions, and second, estimating $Z$ by AIS, which is better suited for handling non-log-concave distributions. The three terms in Eq. 15 arise from (i) ensuring the accuracy of $\widehat{Z}_{0}$, (ii) controlling $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$, and (iii) controlling $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$, respectively, as discussed in 2. above. Due to the non-log-concavity of $\pi$, the action $\mathcal{A}$ is typically large, making (iii), the cost for controlling the discretization error, the dominant complexity. Finally, the $\varepsilon$-dependence can be interpreted as the total duration $T=\Theta\left(\frac{1}{\varepsilon^{2}}\right)$ required for the continuous dynamics to converge (as in Thm. 6) divided by the step size $\widetilde{\Theta}(\varepsilon^{2})$ to control the discretization error.
6 Normalizing Constant Estimation via Reverse Diffusion Sampler
@@
From the analysis of JE and AIS (Thms. 6 and 8), the choice of the interpolation curve $(\pi_{\theta})_{\theta\in[0,1]}$ is crucial for the complexity of AIS. The geometric interpolation (Eq. 11) is widely adopted in practice due to the availability of closed-form scores of the intermediate distributions $\pi_{\theta}$. For certain structured non-log-concave distributions, the associated action is polynomial in the problem parameters, enabling efficient AIS. For instance, Guo et al. (2025, Ex. 2) analyzed a Gaussian mixture target distribution with identical covariance, means having the same norm, and arbitrary weights. However, for general target distributions, the action of the related curve can grow prohibitively large. To illustrate this, we establish an exponential lower bound on the action of a curve starting from a Gaussian mixture, highlighting the potential inefficiency of AIS under geometric interpolation. Our key technical tool is a closed-form expression of the $\text{W}_{\text{2}}$ distance in $\mathbb{R}$ expressed by the inverse cumulative distribution functions (c.d.f.s) of the involved distributions. We then lower bound the metric derivative near the target distribution, where the curve changes the most drastically. The proof of this result is detailed in Sec. E.1.
@@
Consider the Gaussian mixture target distribution $\pi=\frac{1}{2}\operatorname{\mathcal{N}}\left(0,1\right)+\frac{1}{2}\operatorname{\mathcal{N}}\left(m,1\right)$ on $\mathbb{R}$ for some sufficiently large $m\gtrsim 1$, whose potential is $\frac{m^{2}}{2}$-smooth. Under the setting in AIS (Thm. 8), define $\pi_{\theta}(x)\propto\pi(x)\mathrm{e}^{-\frac{\lambda(\theta)}{2}x^{2}}$, $\theta\in[0,1]$, where $\lambda(\theta)=m^{2}(1-\theta)^{r}$ for some $1\leq r\lesssim 1$. Then, the action of the curve $(\pi_{\theta})_{\theta\in[0,1]}$ is lower bounded by $\mathcal{A}_{r}\gtrsim m^{4}\mathrm{e}^{\frac{m^{2}}{40}}$.
Motivated by RDS, we propose leveraging the curve along the OU process in AIS. To support this idea, we first present the following proposition, whose proof is available in Sec. E.2.
@@
Let $\overline{\pi}_{t}$ be the law of $Y_{t}$ in the OU process (Eq. 3) initialized from $Y_{0}\sim\pi\propto\mathrm{e}^{-V}$, where $V$ is $\beta$-smooth and let $m^{2}:=\operatorname{\mathbb{E}}_{\pi}\|\cdot\|^{2}<\infty$. Then, $\int_{0}^{\infty}|\dot{\overline{\pi}}|^{2}_{t}\mathrm{d}t\leq d\beta+m^{2}$.
This proposition shows that under fairly weak conditions on the target distribution, the action of the curve along the OU process, $(\overline{\pi}_{T-t})_{t\in[0,T]}$, behaves much better than Eq. 11. Hence, our analysis of JE (Thm. 6) suggests that this curve is likely to yield more efficient normalizing constant estimation. Furthermore, recall that in our earlier proof, we introduced a compensatory drift term $v_{t}$ to eliminate the lag in ALD. The same principle applies here: ensuring $X_{t}$ precisely following the reference trajectory is advantageous, which results in the time-reversal of OU process (Eq. 4). Building on this insight, we propose an RDS-based algorithm for normalizing constant estimation, and establish a framework for analyzing its oracle complexity. The proof is in Sec. E.3.
@@
Assume a total time duration $T$, an early stopping time $\delta\geq 0$, and discrete time points $0=t_{0}<t_{1}<...<t_{N}=T-\delta\leq T$. For $t\in[0,T-\delta)$, let $t_{-}$ denote $t_{k}$ if $t\in[t_{k},t_{k+1})$. Let $s_{\cdot}\approx\nabla\log\overline{\pi}_{\cdot}$ be a score estimator, and $\phi=\operatorname{\mathcal{N}}\left(0,I\right)$. Consider the following two SDEs on $[0,T-\delta]$ representing the sampling trajectory and the time-reversed OU process, respectively:
@@
$\log\phi(X_{0})+V(X_{T-\delta})+(T-\delta)d+\int_{0}^{T-\delta}\left(\|s_{T-t_{-}}(X_{t_{-}})\|^{2}\mathrm{d}t+\sqrt{2}\left\langle s_{T-t_{-}}(X_{t_{-}}),\mathrm{d}B_{t}\right\rangle\right).$
@@
Then, to ensure $\widehat{Z}$ satisfies Eq. 5, it suffices that $\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$ and $\operatorname{TV}(\pi,\overline{\pi}_{\delta})\lesssim\varepsilon$.
@@
For detailed implementation of this algorithm including the update rule in Eq. 16 and the computation of $W(X)$, see Algorithm 2. To determine the overall complexity, we leverage existing results for RDS (Huang et al., 2024a, b, He et al., 2024, Vacher et al., 2025) to derive the oracle complexity to achieve $\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$. When early stopping is needed (i.e., $\delta>0$), we establish in Lem. 25 that choosing $\delta\asymp\frac{\varepsilon^{2}}{\beta^{2}d^{2}}$ suffices to ensure $\varepsilon$-closeness in TV distance between $\overline{\pi}_{\delta}$ and $\pi$, under weak assumptions similar to Assump. 2. The detailed complexity analysis is deferred to Sec. E.5.
@@
As discussed, RDS can be viewed as an optimally compensated ALD using the OU process as the trajectory. We conclude this section by contrasting these two approaches. On the one hand, analytically-tractable curves such as the geometric interpolation offer closed-form drift terms at all time points, but may exhibit poor action properties (Prop. 9) or bad isoperimetric constants (Chehab et al., 2025), making annealed sampling challenging. On the other hand, alternative curves like the OU process may have better properties in action and isoperimetric constants, but their drift terms, often related to the scores of the intermediate distributions, lack closed-form expressions, and estimating these terms is also non-trivial. This highlights a fundamental trade-off between the complexity of the drift term estimation and the property of the interpolation curve.
7 Conclusion and Future Work
@@
In this paper, we analyzed the complexity of normalizing constant estimation using JE, AIS, and RDS, establishing non-asymptotic convergence guarantees based on insights from continuous-time analysis. Our analysis of JE (Thm. 6) applies to general interpolation curves without requiring explicit isoperimetric assumptions, which significantly extends prior work limited to log-concave distributions. While our main results (Thms. 6 and 8) provide upper complexity bounds, their tightness remains an open question. Deriving general lower bounds would further clarify whether curves with large action inherently require more oracle calls for both sampling and normalizing constant estimation, thereby rigorously validating the arguments in Sec. 6. We also conjecture that our proof techniques can be further extended to samplers beyond overdamped LD (e.g., Hamiltonian or underdamped LD (Sohl-Dickstein and Culpepper, 2012)), and may be applied to estimating normalizing constants of compactly supported distributions on $\mathbb{R}^{d}$ (e.g., convex bodies volume estimation (Cousins and Vempala, 2018)) and discrete distributions (e.g., Ising model and restricted Boltzmann machines (Huber, 2015, Krause et al., 2020)) via the Poisson stochastic integral framework (Ren et al., 2025a, b). We leave these directions for future research.
Appendix A Preliminaries (Continued)

### v2

@@ 4 Analysis of the Annealed Importance Sampling
(II) Show that it suffices to obtain an accurate estimate $\widehat{Z}_{0}$ and initialization distribution $\widehat{\pi}_{0}$, together with sufficiently small KL divergences $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$ and $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$, which quantify the closeness between the continuous dynamics and the discretization error in implementation, respectively.
@@
(III) Using the chain rule, decompose $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$ into the sum of KL divergences between each pair of transition kernels $F_{\ell}$ and $F^{*}_{\ell}$ (i.e., green “distances”). As in Thm.˜2, $F^{*}_{\ell}$, a transition kernel from $\pi_{\theta_{\ell-1}}$ to $\pi_{\theta_{\ell}}$, is realized by ALD with a compensatory vector field, ensuring the SDE exactly follows the trajectory $(\pi_{\theta})_{\theta\in[\theta_{\ell-1},\theta_{\ell}]}$. Similarly, by applying the chain rule and Girsanov theorem, we can express $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$ as the sum of the blue “distances”, allowing for a similar analysis.
@@
(IV) Finally, derive three necessary conditions on the time steps $\theta_{\ell}$ to control both $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$ and $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$. Choosing a proper schedule yields the desired complexity bound.
@@
Our proposed algorithm consists of two phases: first, estimating $Z_{0}$ by TI, which is provably efficient for well-conditioned distributions, and second, estimating $Z$ by AIS, which is better suited for handling non-log-concave distributions. The three terms in ˜11 arise from (i) ensuring the accuracy of $\widehat{Z}_{0}$, (ii) controlling $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$, and (iii) controlling $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})$, respectively, as discussed in (II) above. Due to the non-log-concavity of $\pi$, the action $\mathcal{A}$ is typically large, making (iii), the cost for controlling the discretization error, the dominant complexity. Finally, the $\varepsilon$-dependence can be interpreted as the total duration $T=\Theta\left(\frac{1}{\varepsilon^{2}}\right)$ required for the continuous dynamics to converge (as in Thm.˜2) divided by the step size $\widetilde{\Theta}(\varepsilon^{2})$ to control the discretization error.
5 Normalizing Constant Estimation via Denoising Diffusion
@@
Disadvantage of Geometric Interpolation. From the analysis of JE and AIS (Thms.˜2 and 4), the choice of the interpolation curve $(\pi_{\theta})_{\theta\in[0,1]}$ is crucial for the complexity. The geometric interpolation ˜7 is widely adopted due to the availability of closed-form scores of the intermediate distributions $\pi_{\theta}$, and for certain structured non-log-concave distributions, the associated action is polynomially large, enabling efficient AIS. For instance, Guo et al. (2025, Ex. 2) analyzed a Gaussian mixture target distribution with identical covariance, means having the same norm, and arbitrary weights. However, for general target distributions, the action of the related curve can grow prohibitively large. We establish an exponential lower bound on the action of a curve starting from a Gaussian mixture, highlighting the potential inefficiency of AIS under geometric interpolation.
@@
Consider the Gaussian mixture target distribution $\pi=\frac{1}{2}\operatorname{\mathcal{N}}\left(0,1\right)+\frac{1}{2}\operatorname{\mathcal{N}}\left(m,1\right)$ on $\mathbb{R}$ for some sufficiently large $m\gtrsim 1$, whose potential is $\frac{m^{2}}{2}$-smooth. Under the setting in AIS (Thm.˜4), define $\pi_{\theta}(x)\propto\pi(x)\mathrm{e}^{-\frac{\lambda(\theta)x^{2}}{2}}$, $\theta\in[0,1]$, where $\lambda(\theta)=m^{2}(1-\theta)^{r}$ for some $r\geq 1$. Then, the action of the curve $(\pi_{\theta})_{\theta\in[0,1]}$, $\mathcal{A}_{r}$, is lower bounded by $\mathcal{A}_{r}\gtrsim m^{4}\mathrm{e}^{\frac{m^{2}}{40}}$.
Full proof is in Sec.˜E.1. The key technical tool is a closed-form expression of the $\text{W}_{\text{2}}$ distance in $\mathbb{R}$ expressed by the inverse cumulative distribution functions (c.d.f.s) of the involved distributions, and we lower bound the metric derivative near the target distribution, where the curve changes the most drastically. This observation provides a novel perspective on the quantitative description of the mass teleportation or mode switching phenomenom (Woodard et al., 2009, Syed et al., 2021, Chemseddine et al., 2025), motivating us to explore alternative curves that can potentially yield smaller action, thereby enhancing the efficiency of normalizing constant estimation.
Reverse Diffusion Samplers. Inspired by score-based generative models (Song et al., 2021), recent advancements have led to the development of multimodal samplers based on reversing the Ornstein-Uhlenbeck (OU) process, such as reverse diffusion Monte Carlo (RDMC, Huang et al. (2024a)), recursive score diffusion-based Monte Carlo (RSDMC, Huang et al. (2024b)), zeroth-order diffusion Monte Carlo (ZODMC, He et al. (2024)), and self-normalized diffusion Monte Carlo [^4] This name is introduced by us as the original paper did not provide a name for the proposed algorithm. (SNDMC, Vacher et al. (2025)). We collectively refer to these methods as the reverse diffusion samplers (RDS). The key idea is to simulate the time reversal of the following OU process, which transforms any target distribution $\pi$ into $\phi:=\operatorname{\mathcal{N}}\left(0,I\right)$ as $T\to\infty$:
$\mathrm{d}Y_{t}=-Y_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~t\in[0,T];~Y_{0}\sim\pi.$ (12)
Let $Y_{t}\sim\overline{\pi}_{t}$. The time-reversal $(Y^{\leftarrow}_{t}:=Y_{T-t}\sim\overline{\pi}_{T-t})_{t\in[0,T]}$ satisfies the SDE
Hence, to draw samples $Y^{\leftarrow}_{0}\sim\pi$, it suffices to approximate the scores $\nabla\log\overline{\pi}_{t}$ and discretize ˜13, which can be implemented in various ways, see Sec.˜E.5 for details.
RDS-based Normalizing Constant Estimation. We now propose to leverage $(\overline{\pi}_{T-t})_{t\in[0,T]}$ in AIS. To support this idea, we first present the following proposition, whose proof is in Sec.˜E.2.
@@
Define $\overline{\pi}_{t}$ as the law of $Y_{t}$ in the OU process ˜12 initialized from $Y_{0}\sim\pi\propto\mathrm{e}^{-V}$, where $V$ is $\beta$-smooth. Let $m^{2}:=\operatorname{\mathbb{E}}_{\pi}\|\cdot\|^{2}<\infty$. Then, $\int_{0}^{\infty}|\dot{}\overline{\pi}|^{2}_{t}\mathrm{d}t\leq d\beta+m^{2}$.
Prop.˜2 shows that under fairly weak conditions on the target distribution, the action of the curve along the OU process, $(\overline{\pi}_{T-t})_{t\in[0,T]}$, behaves much better than ˜7. Hence, our analysis of JE (Thm.˜2) suggests that this curve is likely to yield more efficient normalizing constant estimation. Furthermore, recall that in our earlier proof, we introduced a compensatory drift term $v_{t}$ to eliminate the lag in JE. The same principle applies here: ensuring $X_{t}$ precisely following the reference trajectory is advantageous, which results in the time-reversal of OU process ˜13. Building on this insight, we propose an RDS-based algorithm for normalizing constant estimation, and establish a framework for analyzing its oracle complexity. See Sec.˜E.3 for the proof.
@@
Assume a total time duration $T$, an early stopping time $\delta\geq 0$, and discrete time points $0=t_{0}<t_{1}<...<t_{N}=T-\delta\leq T$. For $t\in[0,T-\delta)$, let $t_{-}$ denote $t_{k}$ if $t\in[t_{k},t_{k+1})$. Let $s_{\cdot}\approx\nabla\log\overline{\pi}_{\cdot}$ be a score estimator, and $\phi=\operatorname{\mathcal{N}}\left(0,I\right)$. Consider the following two SDEs on $0,T-\delta]$ representing the sampling trajectory and the time-reversed OU process, respectively:
@@
$\log\phi(X_{0})+V(X_{T-\delta})-(T-\delta)d+\int_{0}^{T-\delta}\left(\|s_{T-t_{-}}(X_{t_{-}})\|^{2}\mathrm{d}t+\sqrt{2}\left\langle s_{T-t_{-}}(X_{t_{-}}),\mathrm{d}B_{t}\right\rangle\right).$
@@
Then, to ensure $\widehat{Z}$ satisfies ˜1, it suffices that $\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$ and $\operatorname{TV}(\pi,\overline{\pi}_{\delta})\lesssim\varepsilon$.
@@
For detailed implementation of the update rule in ˜14 and the computation of $W(X)$, see Algorithm˜2. To determine the overall complexity, we can leverage existing results for RDS to derive the oracle complexity to achieve $\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$. When early stopping is needed (i.e., $\delta>0$), we prove in Lem.˜8 that choosing $\delta\asymp\frac{\varepsilon^{2}}{\beta^{2}d^{2}}$ suffices to ensure $\varepsilon$-closeness in TV distance between $\overline{\pi}_{\delta}$ and $\pi$, under weak assumptions similar to Assump.˜2. For RDMC, RSDMC, ZODMC, and SNDMC, the total complexities are, respectively, $O\left(\operatorname{poly}\left(d,\frac{1}{\zeta}\right)\exp\left(\frac{1}{\varepsilon}\right)^{O(n)}\right)$, $\exp\left(\beta^{3}\log^{3}\left(\beta,d,m^{2},\frac{1}{\zeta}\right)\right)$, $\exp\left(\widetilde{O}(d)\log\beta\log\frac{1}{\varepsilon}\right)$, and $O\left(\left(\frac{\beta(m^{2}\vee d)}{\varepsilon}\right)^{O(d)}\right)$. The full analysis can be found in Sec.˜E.5.
@@
As discussed, RDS can be viewed as an optimally compensated ALD using the OU process as the trajectory. We conclude this section by contrasting these two approaches. On the one hand, analytically-tractable curves such as the geometric interpolation offer closed-form drift terms at all time points, but may exhibit poor action properties (Prop.˜1) or bad isoperimetric constants (Chehab et al., 2025), making annealed sampling challenging. On the other hand, alternative curves like the OU process may have better properties in action and isoperimetric constants, but their drift terms, often related to the scores of the intermediate distributions, lack closed-form expressions, and estimating these terms is also non-trivial. This highlights a fundamental trade-off between the complexity of the drift term estimation and the property of the interpolation curve.
Experiments. We now compare the performance of the methods of normalizing constant estimation for non-log-concave distributions that have been discussed in the paper, including TI, AIS, and the four RDS-based methods. We consider two multimodal target distributions: a modified Müller Brown (MMB) distribution and Gaussian mixture (GM) with $4$ components, both in $\mathbb{R}^{2}$. The quantitative results are summarized in Tab.˜1, where we report the relative error of $\widehat{Z}$ and, for GM, the maximum mean discrepancy (MMD) and $\text{W}_{\text{2}}$ distance between the generated samples $\widehat{\pi}_{\mathrm{samp}}$ and ground truth samples from $\pi$. All RDS-based methods provide accurate estimates of the normalizing constant and high quality samples, while TI and AIS (based on geometric annealing) produce seriously biased estimate due to lack of mode coverage. Further details are presented in App.˜I.
Table 1: Quantitative results of normalizing constant estimation (mean $\pm$ std), best in bold.
Target Metric TI AIS RDMC RSDMC ZODMC SNDMC
MMB $\widehat{Z}/Z$ $0.7527\pm 0.0086$ $2.9740\pm 7.6705$ $0.9829\pm 0.2116$ $1.2885\pm 12.7843$ $0.9878\pm 0.1154$ $\mathbf{1.0053\pm 0.1192}$
GM $\widehat{Z}/Z$ $0.2427\pm 0.0016$ $0.2042\pm 0.0008$ $\mathbf{1.0001\pm 0.0850}$ $0.9202\pm 1.0276$ $0.9766\pm 0.2835$ $0.9973\pm 0.0834$
$\mathrm{MMD}(\widehat{\pi}_{\mathrm{samp}},\pi)$ $2.5407\pm 0.0281$ $2.4618\pm 0.0270$ $0.3581\pm 0.0366$ $0.3124\pm 0.0395$ $0.2591\pm 0.0381$ $\mathbf{0.1576\pm 0.0279}$
$W_{2}(\widehat{\pi}_{\mathrm{samp}},\pi)$ $10.5602\pm 0.0794$ $10.4842\pm 0.0851$ $7.0242\pm 0.9104$ $2.6012\pm 0.2482$ $2.4506\pm 0.2963$ $\mathbf{1.5494\pm 0.6820}$
6 Conclusion, Limitations, and Future Directions
@@
This paper investigates the complexity of normalizing constant estimation using JE, AIS, and RDS, and takes a first step in establishing non-asymptotic convergence guarantees based on insights from continuous-time analysis. Our analysis of JE (Thm.˜2) applies to general interpolations without explicit dependence of isoperimetry, thereby substantially extending prior work limited to log-concave distributions. Several limitations remain: the tightness of our upper bounds (Thms.˜2 and 4) are unknown; the lower bound on the action in Prop.˜1 does not directly imply that JE needs exponentially long time to converge; though the action provides a clean analysis of the statistical efficiency of annealing–which isoperimetric inequalities cannot deal with–its practical interpretability is not well understood. Finally, we conjecture that our proof techniques can extend to samplers beyond overdamped LD (e.g., Hamiltonian or underdamped LD (Sohl-Dickstein and Culpepper, 2012)), and may also apply to estimating normalizing constants of compactly supported distributions on $\mathbb{R}^{d}$ (e.g., convex bodies volume estimation (Cousins and Vempala, 2018)) or discrete distributions (e.g., Ising model and restricted Boltzmann machines (Huber, 2015, Krause et al., 2020)) via the Poisson stochastic integral framework (Ren et al., 2025a, b), which we leave as a direction for future research.
Appendix A Preliminaries
[…]
A.1 Stochastic Analysis: Forward-backward SDEs and Girsanov Theorem
For a stochastic differential equation (SDE) $X=(X_{t})_{t\in[0,T]}$ defined on $\Omega=C([0,T];\mathbb{R}^{d})$, the distribution of $X$ over $\Omega$ is called the path measure of $X$, defined by $\mathbb{P}^{X}$: measurable $A\subset\Omega\mapsto\operatorname{Pr}(X\in A)$. The following lemma, as a corollary of the Girsanov theorem (Üstünel and Zakai, 2013, Prop. 2.3.1 & Cor. 2.3.1), provides a method for computing the Radon-Nikodým (RN) derivative and KL divergence between two path measures, which serves as a key technical tool in our proof.
Lemma 2.
Assume we have the following two SDEs with $t\in[0,T]$:
Remark 3.